The Nutty Shoppe has 10 kg of mixed pecans and cashews, which sell for $ 8 per kilogram. Cashews alone sell for $8 per kilogram, and pecans sell for $9 per kilogram. How many kilograms of pecans are there in the mix?
How many kilograms of cashews are there in the mix?
Thanks

Answer by checkley71(8403)

IF YOU HAVE THE CORRECT PROBLEM THEN 8*10=9P+8C OR 8*10=9*(10-C)+8(C) OR
80=90-9C+8C OR 80-90=-C OR C=10 KG OR CACHEWS THEREFORE THERE ARE 0 KG OF PECANS.